OpenStudy (babyslapmafro):

Could someone please explain how to do the following limit-problem? Evaluate the limit using the substitution z=(x^2)+(y^2) and observing that z approaches 0 from the right if and only if (x,y) approaches (0,0). (click to see limit)

OpenStudy (babyslapmafro):

\[\lim_{(x,y) \rightarrow (0,0)}\frac{ \sin(x^{2}+y ^{2}) }{x^{2}+y ^{2} }\]

OpenStudy (babyslapmafro):

@ganeshie8

ganeshie8 (ganeshie8):

@Jhannybean

OpenStudy (babyslapmafro):

@dan815

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sin x}{x}\]

OpenStudy (anonymous):

=1

OpenStudy (babyslapmafro):

and how did you obtain that limit given the original?

OpenStudy (anonymous):

i didnt, its just something you should remember

OpenStudy (anonymous):

if you're going to make a "z" substitutuion z= x^2 + y^2 then your limit will become \[\lim_{(x,y) \rightarrow (0,0)}\frac{ \sin (z)}{z}\]

OpenStudy (anonymous):

however, since the expression is now in terms of z, you need to convert the limit in relation to z as x-> 0 and y - > 0 we also know x^2 + y^2 = z therefore \[ z = \lim_{(x,y) \rightarrow (0,0)} x^2 + y^2\] (note, this z is the limit, not an actual value that you will use to substitute into the substituted equation) anyways solving this limit will tell us as x and y approaches zero, then z is approaching zero rewriting the limit will give us \[\lim_{z \rightarrow 0} \frac{ \sin (z) }{z}\]

OpenStudy (anonymous):

now what i wrote previously is something that should be memorized, \[\lim_{x \rightarrow 0} \frac{ \sin (x)}{x} = 1\]

OpenStudy (anonymous):

please dont ask me to prove it because i dont remember the proof

OpenStudy (babyslapmafro):

Ok, but using l'hopital's rule works as well, correct?

OpenStudy (babyslapmafro):

Because you have 0/0

OpenStudy (anonymous):

yes it does, but you really should just remember that lim x-> 0 for sin x /x = 1

OpenStudy (babyslapmafro):

the result is then cos(z)/1

OpenStudy (babyslapmafro):

which equals 1

OpenStudy (anonymous):

its a basic identity

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (anonymous):

any questions?

OpenStudy (babyslapmafro):

no i'm ok